# Beam-test evaluation of pre-production Low Gain Avalanche Detectors for the ATLAS High Granularity Timing Detector

### v1

@@ 1 Introduction
The High Granularity Timing Detector (HGTD) [2091129] will be installed in the ATLAS experiment [ATLAS:2008xda] during the so-called Phase-II upgrade for the High-Luminosity period of the LHC (HL-LHC) [Apollinari:2017lan]. The HGTD pixel detector will be introduced in the ATLAS detector at the gap region between the tracker and the end-cap calorimeters at a distance of $\pm$3.5 m from the interaction point covering the pseudorapidity range of 2.4 $|\eta|<$ 4.0. The main goal of this detector is to mitigate the pile-up issue due to the increase in the instantaneous luminosity during the high-luminosity phase of the LHC by combining the time measurements of the hits with the spatial information provided by the Inner Tracker. In addition, the HGTD detector offers unique capabilities for online and offline luminosity determination, an important requirement for precision physics measurements.
The HGTD will consist of 3.16 million silicon-based Low Gain Avalanche Detectors (LGAD) [Pellegrini:2014lki]. LGADs are n-in-p diodes structures with an additional p+ type gain layer, which enhances the rise time and the signal-to-noise ratio, allowing therefore for excellent time resolution. LGAD sensors have been extensively studied, before and after irradiation, in the past years [Allaire:2018bof, Agapopoulou:2020gnf, Lange:2017pxs, Agapopoulou:2022vxk]. Due to the expected high radiation levels, the HGTD is designed to withstand a 1 MeV neutron-equivalent fluence of up to $2.5\times 10^{15}\text{n}_{\mathrm{eq}}/\text{cm}^{2}$ and a total ionizing dose of 2 MGy, estimated at a distance of 120 mm from the beam pipe [2091129]. The most exposed regions are expected to receive radiation damage beyond this level and will therefore require replacement. In order to reach the HGTD design performance, the LGADs must meet targets regarding time resolution of $40\text{\,}\mathrm{ps}$ ($50\text{\,}\mathrm{ps}$) per hit, collected charge $15\text{\,}\mathrm{fC}$ ($4\text{\,}\mathrm{fC}$) and hit efficiencies >> 97% (>> 95%) in the central part of the pad at the start (end) of their lifetime. These requirements should be achieved taking into account a discriminator threshold of about $2\text{\,}\mathrm{fC}$ of the ALTIROC, front-end ASIC foreseen for HGTD [deLaTaille:2018jyb].
@@ 2 Sensor
To study the LGAD performance after irradiation, the sensors were exposed to fluences up to $2.5\times 10^{15}~\text{n}_{\mathrm{eq}}/\text{cm}^{2}$ at the TRIGA reactor [Snoj:2012dib, Ambrozic:2017] in Ljubljana, Slovenia with fast neutrons. Table 1 lists the LGAD sensors measured in the test beams, including the vendor, physical dimensions, and the irradiation fluence. In the table the first column includes the device name assigned to each sensor for easier reference in the text: the sensor name is a concatenation of the sensor design (IHEP or USTC), the wafer number and the fluence. For instance, USTC-W7-2.5 refers to a USTC LGAD from wafer 7, exposed to a fluence of $2.5\times 10^{15}~\text{n}_{\mathrm{eq}}/\text{cm}^{2}$. In one case, to distinguish two sensors, the position of the sensor in the wafer has been added in parenthesis.
@@ 3 Test beam set-up
The measurement of the HGTD pre-production LGADs were conducted during two beam-test campaigns at CERN SPS [SPS_H6] H6A beam line using a high-momentum $120\text{\,}\mathrm{GeV}$ pion beam and one campaign at the DESY II test beam facility [Diener:2018qap] using a $5\text{\,}\mathrm{GeV}$ electron beam. The data-taking set-ups at CERN and DESY were similar, with the main difference being the cooling systems used for the devices under test (DUTs). At CERN, a chiller was employed, ensuring stable operation at -30C, whereas at DESY cooling was provided by dry ice packs, and measurements were performed in the range of -40C to -25C. Figure 3 shows the set-up at CERN and at DESY.
@@
The LGADs were mounted on custom readout boards developed at the University of Santa Cruz  [Agapopoulou:2022vxk, Galloway:2017gfx] with an on-board amplification stage, as shown in Figure 4. The first internal amplifier was followed by an external commercial second-stage 20 dB amplifier to give a trans-impedance of 4700 $\Omega$. The second stage amplifier was placed either inside the cooling box or outside at room temperature. To account for variations in results due to changes in the amplifier trans-impedance with temperature, a 10% systematic uncertainty was assigned, based on a comparison of benchmark measurements with the second-stage amplifier inside versus outside the cooling box.
@@
The time resolution of each DUT was estimated using a Micro-Channel Plate Photomultiplier Tube (MCP-PMT) HPK R3809U-50 [Hamamatsu:R3809U-50] as a time reference. The MCP-PMT was calibrated using coincidences between the MCP-PMT and two DUTs. The time resolution was extracted by considering only the overlapping areas, yielding a measured value of 10.6 $\pm$ 2.2 ps at an operating voltage of 2650 V. The results were cross-checked in dedicated runs at SPS using two additional MCP-PMTs in coincidence, which gave a measured resolution of about 13 ps, in agreement with the previous measurement within uncertainties.
An eight-channel oscilloscope was used to sample the waveforms from both the DUTs and the MCP-PMT, as shown in Figure 5. The oscilloscope has a 1 GHz bandwidth, a 6.25 GS/s sample rate with 8-bit resolution and a memory depth allowing the acquisition of up to 50k samples per channel per trigger cycle. The setup also included a EUDET-type telescope based on six MIMOSA [Jansen:2016bkd] pixel planes (three placed on each side of the DUT) to reconstruct the trajectory of the incident charged particles. The telescope provided the hit position of each track in the DUT frame, enabling efficiency and timing studies.
Triggering was provided by a FE-I4 [Benoit_2016] readout chip–based module, which initiated the acquisition of both oscilloscope waveforms and telescope data. The FE-I4 was configured to accept signals only in a region of interest (ROI) matching the DUT geometries. Trigger logic and distribution were managed by a programmable Trigger Logic Unit (TLU) [Baesso:2019smg]. Busy signals were propagated within the system: in particular, a waveform generator was used to rise a busy signal during the oscilloscope readout, while the TLU also distributed busy states from the telescope and FE-I4. In this way, triggers were vetoed during dead times, ensuring synchronized and loss-free data collection. In addition, the FE-I4 signal was recorded and subsequently used in the offline analysis.
The entire setup was integrated under a DAQ framework based on the EUDAQ2 software [Liu:2686241], which ensured synchronization of the TLU, telescope, and FE-I4 oscilloscope data streams. In addition, the HV power supply of the DUTs was controlled through EUDAQ2, enabling unsupervised and automatic data-taking over different voltage configurations. Figure 5 summarises the DAQ system workflow implemented in the test-beam campaigns.
@@ 4.2 Track reconstruction and data analysis
The reconstruction of particle tracks and subsequent data analysis were performed using the Corryvreckan framework [Dannheim_2021], a modular software package optimized for test beam studies. Data from the six-plane MIMOSA telescope, the FE-I4 reference plane were processed through a standardized Corryvreckan pipeline.
@@ 5.1 Collected Charge
Figure 6: Charge distribution for the sensor IHEP-W16-1.5 operated at 390V. The distribution was fitted with a Landau-Gaussian convoluted function. The collected charge, defined as the most probable value, is 7.36 $\pm$ 0.02 fC. The quoted uncertainty reflects only the statistical component.
@@
Figure 7: Collected charge as a function of bias voltage for IHEP (a) and USTC (b) sensors at various irradiation fluences. The blue (black) dashed line indicates the minimum required charge of 15 fC (4 fC) for unirradiated (irradiated) sensors.
@@ 5.2 Collected Charge as a function of the incident angle
The LGAD performance was studied as a function of the angle of the incident beam. Figure 8 shows the collected charge as a function of the beam incident angle for an unirradiated IHEP sensor and an irradiated USTC sensor. In both cases, the collected charge increases with the incident angle. This effect is due to electrons drifting towards the gain layer along the projection of the particle track. When the beam is perpendicular to the sensor (0 degrees), the charge density on the gain layer is highly localized and there is a potential charge screening [Jimenez-Ramos:2022]. Differently, angled tracks produce a more dispersed charge and thus a reduced charge density across the gain layer, making the suppression mechanism less effective. In addition, the collected charge increases also because a longer particle path length through the sensor results in a larger signal amplitude, and therefore a larger collected charge. This effect increases the collected charge by a factor $1/\cos(\alpha)$ where $\alpha$ is the DUT tilted angle. The increase in collected charge with incident angle is even more apparent when examining the relative difference (in %) of the collected charge with respect to the 0 angle, as shown in Figure 9. Both the screening effect and the particle path inside the sensor are larger for larger value of collected charge, thus the relative increase of charge (MPV) is more significant for unirradiated sensors at higher bias voltage. For irradiated sensors, the relative difference follows the expected trend (magenta line) due to the longer path, demonstrating that the screening effect becomes negligible.
Figure 8: Collected charge as a function of the angles for IHEP-W10-0 (a) and USTC-W15(P49)-2.5 (b) sensors at various bias voltage points.
Figure 9: Relative difference (in %) of the collected charge with respect to angle 0 for IHEP-W10-0 (a) and USTC-W15(P49)-2.5 (b) sensors at various bias voltage points. The magenta line represents the expected charge increased due to longer path of the charge inside the sensor.
@@ 5.3 Time resolution
Figure 10: TOA difference between the sensor IHEP-W16-1.5, operated at 390V, and the MCP-PMT. The distribution is fitted with a Gaussian function, and the width $\sigma$ is extracted from the fit parameters. The quoted uncertainty reflects only the statistical component.
Figures 11 and  12 show the time resolution as a function of the bias voltage for the USTC and IHEP sensors, respectively. Due to differences in the bias voltage settings for sensors with different irradiation levels, the plots are separated into unirradiated and irradiated cases. As before, the error bars represent the contribution from the systematic uncertainty assigned to the experimental setup. For the least irradiated USTC sensors (yellow and brown points on Fig. 11(b)) and the unirradiated IHEP sensors (grey points on Fig. 12(a)) the time resolution worsens at higher bias voltages. This degradation is attributed to noise hits, which become significant when operating the sensors at bias voltages close to the breakdown voltage. All sensors meet the requirement of a time resolution below 40 ps for unirradiated sensors and below 50 ps for irradiated sensors.
Figure 11: Time resolution as a function of the bias voltage for USTC unirradiated (a) and irradiated (b) sensors. The blue (black) dashed line at 40 ps (50 ps) represents the sensor-only per-hit time resolution requirements for unirradiated (irradiated) sensors.
Figure 12: Time resolution as a function of the bias voltage for IHEP unirradiated (a) and irradiated (b) sensors. The blue (black) dashed line at 40 ps (50 ps) represents the sensor-only per-hit time resolution requirements for unirradiated (irradiated) sensors.
@@ 5.4 Time resolution as a function of the incident angle
Figure 13: Time resolution as a function of the angles for IHEP-W10-0 (a) and USTC-W15(P49)-2.5 (b) sensors at various bias voltage points.
Figure 14: Relative difference (in %) of the time resolution with respect to angle 0 for IHEP-W10-0 (a) and USTC-W15(P49)-2.5 (b) sensors at various bias voltage points.
@@ 5.5 Hit Reconstruction Efficiency
where $Q_{\text{cut}}$ is the minimum collected charge, set to 2 fC. The fiducial region is defined as the rectangle centered 60% area of the DUT [^6] The central area is evaluated considering -0.81 (-0.87) mm ¡ x ¡ 0.81 (0.87) mm and -0.31 (-0.29) mm ¡ y ¡ 0.31 (0.29) mm for USTC (IHEP) sensors., excluding edge effects, where the sensor response is expected to be uniform. The hit reconstruction efficiency is calculated only for events with exactly one reconstructed track to discard possible pile-up events. The efficiency study has been performed only using data collected at CERN.
@@
Figure 15: Hit reconstruction efficiency versus bias voltage for sensors irradiated at $2.5\times 10^{15}~n_{\text{eq}}/\text{cm}^{2}$. Results are shown for multiple sensor designs from USTC and IHEP. The dashed line indicates the 95% efficiency requirement for HGTD operation. A charge threshold of $Q_{\text{cut}}>2~\mathrm{fC}$ is applied.
@@
Figure 16: 2D maps of the efficiency as a function of the hit position in the sensor plane for IHEP-W16-0 sensor operated at a bias voltage of 125 V (a) and for the USTC-W15(P1)-2.5 sensor operated at a bias voltage of 555 V (c).
@@
Figure 17: Projections on the x- and y-axis of the efficiency in the central region of the sensor for USTC (a,b) and IHEP (c,d) sensors. Each plot shows the comparison of the projection of the efficiency for unirradiated (red line) and irradiated sensors (blue line). For USTC-W7-0, the right tail of the distribution is truncated to avoid spikes due to the missing statistics. The projections on the x-axis (y-axis) are evaluated considering the integral of the central region from -0.31 mm to 0.31 mm (from -0.81 mm to 0.81 mm) and from -0.29 mm to 0.29 mm (from -0.87 mm to 0.87 mm) for USTC and IHEP sensors, respectively.

### v2

@@ 1 Introduction
The High Granularity Timing Detector (HGTD) [12] will be installed in the ATLAS experiment [7] during the so-called Phase-II upgrade for the High-Luminosity period of the LHC (HL-LHC) [6]. The HGTD pixel detector will be introduced in the ATLAS detector at the gap region between the tracker and the end-cap calorimeters at a distance of $\pm$3.5 m from the interaction point covering the pseudorapidity range of 2.4 $|\eta|<$ 4.0. The main goal of this detector is to mitigate the pile-up issue due to the increase in the instantaneous luminosity during the high-luminosity phase of the LHC by combining the time measurements of the hits with the spatial information provided by the Inner Tracker. In addition, the HGTD detector offers unique capabilities for online and offline luminosity determination, an important requirement for precision physics measurements.
The HGTD will consist of 3.16 million silicon-based Low Gain Avalanche Detectors (LGAD) [28]. LGADs are n-in-p diodes structures with an additional p+ type gain layer, which enhances the rise time and the signal-to-noise ratio, allowing therefore for excellent time resolution. LGAD sensors have been extensively studied, before and after irradiation, in the past years [4, 1, 23, 2]. Due to the expected high radiation levels, the HGTD is designed to withstand a 1 MeV neutron-equivalent fluence of up to $2.5\times 10^{15}\text{n}_{\mathrm{eq}}/\text{cm}^{2}$ and a total ionizing dose of 2 MGy, estimated at a distance of 120 mm from the beam pipe [12]. The most exposed regions are expected to receive radiation damage beyond this level and will therefore require replacement. In order to reach the HGTD design performance, the LGADs must meet targets regarding time resolution of $40\text{\,}\mathrm{ps}$ ($50\text{\,}\mathrm{ps}$) per hit, collected charge $15\text{\,}\mathrm{fC}$ ($4\text{\,}\mathrm{fC}$) and hit efficiencies >> 97% (>> 95%) in the central part of the pad at the start (end) of their lifetime. These requirements should be achieved taking into account a discriminator threshold of about $2\text{\,}\mathrm{fC}$ of the ALTIROC, front-end ASIC foreseen for HGTD [14].
@@ 2 Sensor
To study the LGAD performance after irradiation, the sensors were exposed to fluences up to $2.5\times 10^{15}~\text{n}_{\mathrm{eq}}/\text{cm}^{2}$ at the TRIGA reactor [30, 5] in Ljubljana, Slovenia with fast neutrons. Table 1 lists the LGAD sensors measured in the test beams, including the vendor, physical dimensions, and the irradiation fluence. In the table the first column includes the device name assigned to each sensor for easier reference in the text: the sensor name is a concatenation of the sensor design (IHEP or USTC), the wafer number and the fluence. For instance, USTC-W7-2.5 refers to a USTC LGAD from wafer 7, exposed to a fluence of $2.5\times 10^{15}~\text{n}_{\mathrm{eq}}/\text{cm}^{2}$. In one case, to distinguish two sensors, the position of the sensor in the wafer has been added in parenthesis.
@@ 3 Test beam set-up
The measurement of the HGTD pre-production LGADs were conducted during two beam-test campaigns at CERN SPS [9] H6A beam line using a high-momentum $120\text{\,}\mathrm{GeV}$ pion beam and one campaign at the DESY II test beam facility [15] using a $5\text{\,}\mathrm{GeV}$ electron beam. The data-taking set-ups at CERN and DESY were similar, with the main difference being the cooling systems used for the devices under test (DUTs). At CERN, a chiller was employed, ensuring stable operation at -30C, whereas at DESY cooling was provided by dry ice packs, and measurements were performed in the range of -40C to -25C. Figure 3 shows the set-up at CERN and at DESY.
@@
The LGADs were mounted on custom readout boards developed at the University of Santa Cruz  [2, 17] with an on-board amplification stage, as shown in Figure 4. The first internal amplifier was followed by an external commercial second-stage 20 dB amplifier to give a trans-impedance of 4700 $\Omega$. The second stage amplifier was placed either inside the cooling box or outside at room temperature. To account for variations in results due to changes in the amplifier trans-impedance with temperature, a 10% systematic uncertainty was assigned to the collected charge and time reconstruction. This uncertainty is based on a comparison of benchmark measurements with the second-stage amplifier inside versus outside the cooling box.
@@
The time resolution of each DUT was estimated using a Micro-Channel Plate Photomultiplier Tube (MCP-PMT) HPK R3809U-50 [22] as a time reference. The MCP-PMT was calibrated using coincidences between the MCP-PMT and two DUTs. The time resolution was extracted by considering only the overlapping areas, yielding a measured value of 10.6 $\pm$ 2.2 ps at an operating voltage of 2650 V. The results were cross-checked in dedicated runs at SPS using two additional MCP-PMTs in coincidence, which gave a measured resolution of about 13 ps, in agreement with the previous measurement within uncertainties.
An eight-channel oscilloscope was used to sample the waveforms from both the DUTs and the MCP-PMT, as shown in Figure 5. The oscilloscope has a 1 GHz bandwidth, a 6.25 GS/s sample rate with 8-bit resolution and a memory depth allowing the acquisition of up to 50k samples per channel per trigger cycle. The setup also included a EUDET-type telescope based on six MIMOSA [20] pixel planes (three placed on each side of the DUT) to reconstruct the trajectory of the incident charged particles. The telescope provided the hit position of each track in the DUT frame, enabling efficiency and timing studies.
Triggering was provided by a FE-I4 [10] readout chip–based module, which initiated the acquisition of both oscilloscope waveforms and telescope data. The FE-I4 was configured to accept signals only in a region of interest (ROI) matching the DUT geometries. Trigger logic and distribution were managed by a programmable Trigger Logic Unit (TLU) [8]. Busy signals were propagated within the system: in particular, a waveform generator was used to rise a busy signal during the oscilloscope readout, while the TLU also distributed busy states from the telescope and FE-I4. In this way, triggers were vetoed during dead times, ensuring synchronized and loss-free data collection. In addition, the FE-I4 signal was recorded and subsequently used in the offline analysis.
The entire setup was integrated under a DAQ framework based on the EUDAQ2 software [26], which ensured synchronization of the TLU, telescope, and FE-I4 oscilloscope data streams. In addition, the HV power supply of the DUTs was controlled through EUDAQ2, enabling unsupervised and automatic data-taking over different voltage configurations. Figure 5 summarises the DAQ system workflow implemented in the test-beam campaigns.
@@ 4.2 Track reconstruction and data analysis
The reconstruction of particle tracks and subsequent data analysis were performed using the Corryvreckan framework [13], a modular software package optimized for test beam studies. Data from the six-plane MIMOSA telescope and the FE-I4 reference plane were processed through a standardized Corryvreckan pipeline.
@@ 5.1 Collected Charge
Figure 6: Charge distribution for the sensor IHEP-W16-1.5 operated at 390V. The measurement was performed at a temperature of -30C. The distribution was fitted with a Landau-Gaussian convoluted function. The collected charge, defined as the most probable value, is 7.36 $\pm$ 0.02 fC. The quoted uncertainty reflects only the statistical component.
@@
Figure 7: Collected charge as a function of bias voltage for IHEP (a) and USTC (b) sensors at various irradiation fluences. All measurements were performed at -30C, except for the IHEP-W2-1.5 and IHEP-W2-2.5 sensors tested at DESY, where measurements were carried out in the temperature range from -40C to -25C. The blue (black) dashed line indicates the minimum required charge of 15 fC (4 fC) for unirradiated (irradiated) sensors.
@@ 5.2 Collected Charge as a function of the incident angle
The LGAD performance was studied as a function of the angle of the incident beam. Figure 8 shows the collected charge as a function of the beam incident angle for an unirradiated IHEP sensor and an irradiated USTC sensor. In both cases, the collected charge increases with the incident angle. This effect is due to electrons drifting towards the gain layer along the projection of the particle track. When the beam is perpendicular to the sensor (0 degrees), the charge density on the gain layer is highly localized and there is a potential charge screening [21]. Differently, angled tracks produce a more dispersed charge and thus a reduced charge density across the gain layer, making the suppression mechanism less effective. In addition, the collected charge increases also because a longer particle path length through the sensor results in a larger signal amplitude, and therefore a larger collected charge. This effect increases the collected charge by a factor $1/\cos(\alpha)$ where $\alpha$ is the DUT tilted angle. The increase in collected charge with incident angle is even more apparent when examining the relative difference (in %) of the collected charge with respect to the 0 angle, as shown in Figure 9. Both the screening effect and the particle path inside the sensor are larger for larger value of collected charge, thus the relative increase of charge (MPV) is more significant for unirradiated sensors at higher bias voltage. For irradiated sensors, the relative difference follows the expected trend (magenta line) due to the longer path, demonstrating that the screening effect becomes negligible.
Figure 8: Collected charge as a function of the angles for IHEP-W10-0 (a) and USTC-W15(P49)-2.5 (b) sensors at various bias voltage points. All measurements were performed at a temperature of -30C.
Figure 9: Relative difference (in %) of the collected charge with respect to angle 0 for IHEP-W10-0 (a) and USTC-W15(P49)-2.5 (b) sensors at various bias voltage points. All measurements were performed at a temperature of -30C. The magenta line represents the expected charge increased due to longer path of the charge inside the sensor.
@@ 5.3 Time resolution
Figure 10: TOA difference between the sensor IHEP-W16-1.5, operated at 390V, and the MCP-PMT. The measurement was performed at a temperature of -30C. The distribution is fitted with a Gaussian function, and the width $\sigma$ is extracted from the fit parameters. The quoted uncertainty reflects only the statistical component.
Figures 11 and  12 show the time resolution as a function of the bias voltage for the USTC and IHEP sensors, respectively. Due to differences in the bias voltage settings for sensors with different irradiation levels, the plots are separated into unirradiated and irradiated cases. As before, the error bars represent the contribution from the systematic uncertainty assigned to the experimental setup. For the least irradiated USTC sensors (yellow and brown points on Fig. 11(b)) and the unirradiated IHEP sensors (grey points on Fig. 12(a)) the time resolution worsens at higher bias voltages. This degradation is attributed to noise hits, which become significant when operating the sensors at bias voltages close to the breakdown voltage. Time resolution differences among unirradiated sensors can arise from variations in charge-carrier velocity saturation, which depends on the electric field [29], and thus on the sensor bias. All sensors meet the requirement of a time resolution below 40 ps for unirradiated sensors and below 50 ps for irradiated sensors.
Figure 11: Time resolution as a function of the bias voltage for USTC unirradiated (a) and irradiated (b) sensors. All measurements were performed at a temperature of -30C. The blue (black) dashed line at 40 ps (50 ps) represents the sensor-only per-hit time resolution requirements for unirradiated (irradiated) sensors.
Figure 12: Time resolution as a function of the bias voltage for IHEP unirradiated (a) and irradiated (b) sensors. All measurements were performed at -30C, except for the IHEP-W2-1.5 and IHEP-W2-2.5 sensors tested at DESY, where measurements were carried out in the temperature range from -40C to -25C. The blue (black) dashed line at 40 ps (50 ps) represents the sensor-only per-hit time resolution requirements for unirradiated (irradiated) sensors.
@@ 5.4 Time resolution as a function of the incident angle
Figure 13: Time resolution as a function of the angles for IHEP-W10-0 (a) and USTC-W15(P49)-2.5 (b) sensors at various bias voltage points. All measurements were performed at a temperature of -30C.
Figure 14: Relative difference (in %) of the time resolution with respect to angle 0 for IHEP-W10-0 (a) and USTC-W15(P49)-2.5 (b) sensors at various bias voltage points. All measurements were performed at a temperature of -30C.
@@ 5.5 Hit Reconstruction Efficiency
where $Q_{\text{cut}}$ is the minimum collected charge, set to 2 fC. The fiducial region is defined as a rectangle covering the central 60% of the DUT area [^6] The central area is evaluated considering -0.81 (-0.87) mm ¡ x ¡ 0.81 (0.87) mm and -0.31 (-0.29) mm ¡ y ¡ 0.31 (0.29) mm for USTC (IHEP) sensors., excluding edge effects, where the sensor response is expected to be uniform. The hit reconstruction efficiency is calculated only for events with exactly one reconstructed track to discard possible pile-up events. The efficiency study has been performed only using data collected at CERN.
@@
Figure 15: Hit reconstruction efficiency versus bias voltage for sensors irradiated at $2.5\times 10^{15}~n_{\text{eq}}/\text{cm}^{2}$. All measurements were performed at a temperature of -30C. Results are shown for multiple sensor designs from USTC and IHEP. The dashed line indicates the 95% efficiency requirement for HGTD operation. A charge threshold of $Q_{\text{cut}}>2~\mathrm{fC}$ is applied.
@@
Figure 16: 2D maps of the efficiency as a function of the hit position in the sensor plane for IHEP-W16-0 sensor operated at a bias voltage of 125 V (a) and for the USTC-W15(P1)-2.5 sensor operated at a bias voltage of 555 V (c). All measurements were performed at a temperature of -30C.
@@
Figure 17: Projections on the x- and y-axis of the efficiency in the central region of the sensor for USTC (a,b) and IHEP (c,d) sensors. All measurements were performed at a temperature of -30C. Each plot shows the comparison of the projection of the efficiency for unirradiated (red line) and irradiated sensors (blue line). For USTC-W7-0, the right tail of the distribution is truncated to avoid spikes due to the missing statistics. The projections on the x-axis (y-axis) are evaluated considering the integral of the central region from -0.31 mm to 0.31 mm (from -0.81 mm to 0.81 mm) and from -0.29 mm to 0.29 mm (from -0.87 mm to 0.87 mm) for USTC and IHEP sensors, respectively.